OpenStudy (anonymous):

Does anyone know what 41.293÷50.6 , 40.478÷50.6 and 42.986÷50.6 ? i will REALLY APPRECIATE IT if anyone can figure these out for me pleaseeeeeee and thank you so much if anyone does .

OpenStudy (anonymous):

41.293÷50.6 =.816 40.478÷50.6=.800 42.986÷50.6=.850

OpenStudy (anonymous):

THANK YOU SO MUCH (:

OpenStudy (anonymous):

You are using a computer right now, @AshleyAlvelo12 , so you must have access to a calculator. Why couldn't you calculate these yourself?